# Taxonomy and Analysis of Sensitive User Queries in Generative AI Search

### v1

@@ Abstract
Although there has been a growing interest among industries to integrate generative LLMs into their services, limited experiences and scarcity of resources acts as a barrier in launching and servicing large-scale LLM-based conversational services. In this paper, we share our experiences in developing and operating generative AI models within a national-scale search engine, with a specific focus on the sensitiveness of user queries. We propose a taxonomy for sensitive search queries, outline our approaches, and present a comprehensive analysis report on sensitive queries from actual users.
@@ 1 Introduction
Despite their advantages, successful launch and maintenance of large-scale LLM-based services has been limited to a few organizations including OpenAI, Google, and Microsoft. The main challenge in creating large-scale LLM-based services has been on the scarcity of computational and human resources required for model pretraining and fine-tuning, however, the recent publicly available open-source LLMs Touvron et al. (2023); Jiang et al. (2023) alleviated the challenges associated with model training significantly.
We believe that next significant obstacle lies in the absence of adequate service experience focusing on user behaviors in conversational settings, particularly with regard to safety considerations encompassing both user inquiries and service responses. While a substantial body of previous research focused on the safety of generative model responses Ousidhoum et al. (2021); Wei et al. (2023); Kumar et al. (2023b), there have been limited studies Qi et al. (2021); Kumar et al. (2023a) examining model security and safety in the context of adversarial prompting and backdoor attacks. Moreover, these studies primarily aimed to secure generative models in lab settings rather than to address sensitive user inputs in publicly available services.
To narrow this gap, this paper focuses on user input, particularly in the context of query sensitiveness111We prefer to use this term over safety because the concept of safety may vary across cultures and service purposes., within a generative search service provided by a leading search portal in Korea. By examining a wide range of sensitive query types, a comprehensive taxonomy of sensitive user queries is proposed. We also share the distribution and details of the sensitive queries from actual users.
By doing so, we aim to offer valuable insights into how users interact with and potentially exploit the system, which is a crucial consideration point in developing and operating such services. We anticipate that this research will aid in creation of other generative services effectively handling sensitive user inputs. It will serve as a reference point for future endeavors, helping to approximate the requirements necessary for end-user services.
@@
We share our experience of designing the input part of a generative LLM service based on a national-scale search engine in Korea.
We (1) introduce a taxonomy for sensitive queries, (2) analyze the distribution of sensitive queries and how they respond to social issues, and (3) provide a detailed keyword analysis for more insights.
@@ 2 Related Works
Generative Applicatiosn
As we mentioned in Section 1, our service shares similarities with other platforms such as chatGPT, Google Bard, and Bing. Notably, Google recently introduced a generative model called Gemini222https://labs.google.com/search/, which is aligned with a search engine similar to ours. Additionally, there may be other successful services in different countries. However, this paper aims to contribute by providing a comprehensive account of our experiences. We have shared our sensitive category taxonomy, insights in managing these categories, and the analysis of log distributions from various perspectives.
Releasing Distribution of Log on a National-Scale Service
@@
Given that our service primarily operates in Korean, this work directly relates to SQuARe Lee et al. (2023a), which provides datasets of sensitive questions and acceptable (or non-acceptable) responses in Korean. However, our work differs from previous studies in terms of focusing on the input queries of a search engine in real-world scenarios. Additionally, while SQuARe categorizes sensitive questions into three categories (contentious, ethical, and predictive), we have developed a more comprehensive taxonomy consisting of 12 detailed categories to effectively handle various sensitive intentions behind search queries. This enables us to generate guided responses that explain why certain information cannot be provided.
@@
Apart from language differences, PALMS Solaiman and Dennison (2021) extensively investigates potential sensitive categories in the outputs of generative models. Although their study examines response generation, it serves as a valuable reference for defining sensitiveness categories in an application, including input queries. Furthermore, we acknowledge that sensitive categories may vary across cultures, and our paper contributes to demonstrating the effectiveness of our taxonomy at the input level and in different cultural contexts.
@@
Llama Guard Inan et al. (2023) illustrates how they constructed a safety model for input prompts and model responses, focusing on safety taxonomy, data collection, and training method. In comparison to our categories, Llama Guard employs six categories: Violence & Hate, Sexual Content, Criminal Planning, Guns & Illegal Weapons, Regulated or Controlled Substances, and Suicide & Self-Harm, assuming human and AI interaction in a conversational context. In comparison, our work considers more extensive usage scenarios within a general search service, covering a wider range of comprehensive areas. Some examples include age-restricted content, copyright infringement, personification of the system, future prediction, and error-inducing queries, among others.
3 Search-based Generative Application
Search Engine.
@@ 3 Search-based Generative Application
Our search portal, which primarily focus in Korea, incorporates email services, blogs, news, and user communities. The search engine that is employed utilizes sophisticated databases and lexical/neural matching to find relevant web pages. As of August 2023, the search service holds a share of 68.9% in Korea333https://www.ajunews.com/view/20230821111740686.
Generative Application Flow.
@@
The generative AI search is designed to fulfill user search intent. As illustrated in Figure 1, it works alongside the search engine in the following way: the system imitates human reasoning to establish a search strategy. The search engine retrieves web pages based on the search keywords. The evidence selector identifies documents with potential answers. Summarization is performed on the selected documents, providing context for response. Factual consistency checks ensure reliable results. Finally, the system provides multifaceted answers to user questions.
@@
Our sensitive query classifier, positioned between Question and Reasoning, is responsible for identifying whether a search query should be treated as sensitive. When a query is flagged as sensitive, the downstream modules utilize this information as a cue to generate a response that prioritizes safety and appropriateness.
@@ 4 Taxonomy of Sensitive Queries
Query sensitiveness encompasses the harmfulness of the query itself as well as the degree of social, economical, and political impacts from the potential service response. Although the definition of sensitiveness may vary depending on purposes of the service and culture of the service area, we believe that it is worthy to have a comprehensive coverage embracing those variations. To this end, we started from reviewing the leading LLM-based service providers’ set of guidelines, including OpenAI444https://openai.com/policies/usage-policies, Google555https://policies.google.com/terms/generative-ai/use-policy?hl=en and Meta666https://ai.meta.com/llama/use-policy/, as well as previous works on the AI safety area Inan et al. (2023) including works considering Korean culture Lee et al. (2023a, b). We then revised it to fit to the purpose and the best user experience of the search engine.
We present three high-level sensitive areas, categorized by the nature of potential issues of the query and response, from the corporate service and social responsibility standpoints: (1) Legal, (2) Ethical, and (3) Service-sensitive issues.
@@ 4.2 Ethical issues
Personification of the system Queries assuming the system as a human, and/or asking the system to perform tasks beyond its capabilities such as "Act like my boy-/girl-friend" and "Could you climb?". This category falls into a sensitive area depending on the purpose and definition of the service. However, queries regarding the pre-defined capabilities of the system (e.g., write an essay) do not have to be considered sensitive.
@@ 4.3 Service-sensitive issues
Error-inducing Instances of queries that elicit inaccurate or unexpected responses, such as hallucination and prompt injection, need to be addressed. For instance, media reports have highlighted cases where LLMs have provided affirmative responses to nonsensical questions, such as "Tell me the date when Samsung launched the latest iPhone?" Additionally, prompt injection attacks pose a risk of exposing confidential training methodologies and datasets, and therefore require careful handling.
@@ 5 Analysis
We analyzed all user queries collected from September 20th, 2023 through November 28th, 2023. This period accounts for a total of 70 days since the launch of our service (Figure 3). While we are unable to disclose the exact numbers of active users and total queries due to confidentiality reasons, we provide a relative overview of the demographic distribution and the daily query volume in relation to the maximum number of daily queries.
To effectively analyze the large volume of user queries, we developed a sensitive query classifier model, which is further detailed in Section 5.1. This model enables us to examine the distribution of sensitive queries as discussed in Section 5.3. We use the query sensitiveness taxonomy proposed in Section 4 for the distribution analysis. We also investigate how the distribution of sensitive queries changes in response to specific social issues that arose in news and social media.
5.1 Sensitive Query Classifier Model
Our model is developed based on HyperCLOVA X backbone model Yoo et al. (2024), whose extension is the state-of-the-art Korean LLM. The training dataset consists of publicly available Korean datasets Moon et al. (2020) as well as internally generated datasets by our red team. To ensure the reliability of our service, we have developed an operation tool for rule-based adjustment so that we can continuously improve the service quality and build feedback loops to refresh the model with the adjustment history records. For more comprehensive understanding of the model, modules and the performance, we provide details in Appendix A.1.

### v2

@@ Abstract
Although there has been a growing interest among industries in integrating generative LLMs into their services, limited experience and scarcity of resources act as a barrier in launching and servicing large-scale LLM-based services. In this paper, we share our experiences in developing and operating generative AI models within a national-scale search engine, with a specific focus on the sensitiveness of user queries. We propose a taxonomy for sensitive search queries, outline our approaches, and present a comprehensive analysis report on sensitive queries from actual users. We believe that our experiences in launching generative AI search systems can contribute to reducing the barrier in building generative LLM-based services.
Taxonomy and Analysis of Sensitive User Queries
in Generative AI Search System
Hwiyeol Jo1,2, Taiwoo Park1††thanks: Now at Google†, Hyunwoo Lee1,2, Nayoung Choi3††thanks: Work carried out at NAVER, Changbong Kim2, Ohjoon Kwon2, Donghyeon Jeon2, Eui-Hyeon Lee2, Kyoungho Shin2, Sun Suk Lim2, Kyungmi KIM2, Jihye Lee2, Sun Kim1††thanks: Co-corresponding author 1NAVER Search US, 2NAVER, 3Emory University {hwiyeol.jo,taiwoo.park,hanu.lee,nayoung.choi,changbong.kim, ohjoon.kwon,donghyeon.jeon,euihyeon.lee,kyoungho.shin, dongle.75,kyungmi.kim,ideal14.jhlee,sunkim.us}@navercorp.com
@@ 1 Introduction
Despite their advantages, the successful launch and maintenance of large-scale LLM-based services has been limited to a few organizations. The main challenge in creating large-scale LLM-based services has been the scarcity of computational and human resources required for model pretraining and fine-tuning, however, the recent publicly available open-source LLMs Touvron et al. (2023); Jiang et al. (2023) alleviated the challenges associated with model training significantly.
We believe that next significant obstacle lies in the absence of adequate service experience focusing on user behaviors in conversational settings, particularly with regard to safety considerations encompassing both user inquiries and service responses. While a substantial body of previous research focused on the safety of generative model responses Ousidhoum et al. (2021); Wei et al. (2023); Kumar et al. (2023b), these studies primarily aimed to secure generative models in lab settings rather than to address sensitive user inputs in publicly available services.
To narrow this gap, this paper focuses on user input, particularly in the context of query sensitiveness111We prefer to use this term over safety because the concept of safety may vary across cultures and service purposes., within a generative search service provided by a leading search portal in Korea. By examining a wide range of sensitive query types, a comprehensive taxonomy of sensitive user queries and considerations behind the taxonomy are proposed. We also share the distribution and details of sensitive queries from actual users.
By doing so, we aim to provide valuable insight into how users interact with and potentially exploit the system, which is a crucial consideration point in developing and operating such services. We anticipate that this research will aid the creation of other generative services that effectively handle sensitive user input. It will serve as a reference point for future endeavors, helping to approximate the requirements necessary for end-user services.
@@
We share our experience in designing the input part of a generative LLM service based on a national-scale search engine in South Korea, from the safety standpoint.
We (1) introduce a taxonomy of sensitive queries for our real-world system, (2) analyze the distribution of sensitive queries and how they respond to social issues, and (3) provide a detailed keyword analysis for more insights.
Lastly, we outline key considerations for constructing the taxonomy and implementations, encompassing ethical aspects like the well-being of annotators and system’s reliability.
@@ 2 Related Works
Generative Model Applications
Our service shares similarities with generative model platforms. Notably, Google introduced a generative model called Gemini222https://labs.google.com/search/, which is aligned with a search engine. In addition, there may be other successful services. In this paper, we aim to contribute by providing a comprehensive account of our experiences. We share our sensitive category taxonomy, insights behind them, and the analysis of log distributions from various perspectives.
Distribution of Logs on a National-Scale Service
@@
Given that our service primarily operates in Korean, this work directly relates to SQuARe Lee et al. (2023a), which provides datasets of sensitive questions and acceptable (or not acceptable) responses in Korean. However, our work differs in terms of focusing on the input queries of a search engine in real-world scenarios. Additionally, while SQuARe categorized sensitive questions into three categories (contentious, ethical, and predictive), we have developed a more comprehensive taxonomy consisting of 12 detailed categories to effectively handle various sensitive intentions behind search queries.
@@
Apart from language differences, PALMS Solaiman and Dennison (2021) extensively investigated potential sensitive categories in the outputs of generative models. Although their study examined response generation, it served as a valuable reference for defining sensitive categories in an application, including input queries. Furthermore, we acknowledge that sensitive categories may vary across cultures, and our paper contributes to demonstrating the effectiveness of our taxonomy at the input level and in a different cultural context.
@@
Llama Guard Inan et al. (2023) illustrates how they constructed a safety model for input prompts and model responses, focusing on safety taxonomy, data collection, and training method. Compared to our categories, Llama Guard employed six categories: Violence & Hate, Sexual Content, Criminal Planning, Guns & Illegal Weapons, Regulated or Controlled Substances, and Suicide & Self-Harm, assuming human and AI interaction in a conversational context. On the other hand, we consider more extensive usage scenarios within a different service type (i.e., search service), covering a wider range. Some examples will be discussed in Section 4, such as age-restricted content, copyright infringement, personification of the system, future prediction, and error-inducing queries, among others.
While we recognize the limitations of our current taxonomy in terms of cultural and service-type coverage, we emphasize that the system’s ongoing operation without significant issues validates its applicability to our target domain. Due to privacy, confidentiality, and potential commercial concerns, we are unable to publicly share complete details such as the exact number of queries and the dataset including system inputs/outputs.
3 Search-based Generative System
Search Engine
@@ 3 Search-based Generative System
Our search portal, focusing primarily in South Korea, incorporates email services, blogs, news, and user communities. Its search engine utilizes sophisticated databases and lexical/semantic matching to find relevant web pages. As of January 2025, the search service holds a share of 63.5% in South Korea333http://www.internettrend.co.kr/trendForward.tsp.
System Flow
@@
As illustrated in Figure 1, it works alongside the search engine in the following way: the system imitates human reasoning to establish a search strategy. The search engine retrieves web pages based on search keywords. The evidence selector identifies documents with potential answers. Summarization is performed on the selected documents, providing context for response. Factual consistency checks ensure reliable results. Finally, the system provides multifaceted answers.
@@
Our sensitive query classifier, positioned between Question and Reasoning, is responsible for identifying whether a search query may be sensitive. When a query is flagged as sensitive, the downstream modules utilize this information as a cue to generate a response that prioritizes safety and appropriateness. The details will be presented in Section 5.2.
System’s Scale
444We understand the importance of providing information about our system’s scale.Due to legal restrictions, we are unable to include specific details in the paper that could potentially reveal the exact numbers. While we cannot disclose the overall system scale, we can offer some publicly available statistics that provide insights.
As of October 2024, our primary search engine portal boasts an impressive 4.8 million daily active users (DAU) on mobile devices. While PC access is also available, specific user statistics are not publicly disclosed. Our portal simplifies the process by providing a simple button that allows users to input search engine queries directly into the generative model. We believe this approach makes generative models accessible to a wider user base.
Furthermore, although exact query numbers are unavailable, we can leverage unofficial keyword analysis tools like keywordsound555https://keywordsound.com/ to estimate keyword-level search volume. For instance, the keyword "넷플릭스 (Netflix)" generates an estimated 1.8 million searches per month, while "누누티비 (Pirate sites offering paid streaming content)" receives approximately 169,000 monthly searches666Note that these numbers are approximations and variations in query phrasing can exist..
@@ 4 Taxonomy of Sensitive Queries
Query sensitiveness encompasses the harmfulness of the query itself as well as the degree of social, economical, and political impacts from the potential service response. Although the definition of sensitiveness may vary depending on purposes of the service and culture of the service area, we believe that it is worthy to have a comprehensive coverage embracing those variations. To this end, we started from reviewing the leading LLM-based service providers’ set of guidelines, including OpenAI777https://openai.com/policies/usage-policies/, Google888https://policies.google.com/terms/generative-ai/use-policy/ and Meta999https://ai.meta.com/llama/use-policy/, as well as previous works on the AI safety area Inan et al. (2023) including works considering Korean culture Lee et al. (2023a, b). We then revised it to fit the service purpose and to provide better user experience of the search engine.
Although our taxonomy is built upon the previous ideas, we strongly believe that the adaptions and modifications for real-world systems are important. Also, operating generative systems with the taxonomy provide an empirical evidence to the research. We finally present three high-level sensitive areas, categorized by the nature of potential issues from the corporate service and social responsibility standpoints: (1) Legal, (2) Ethical, and (3) Service-sensitive issues. Each high-level category consists of detailed areas with a quick description and a set of examples.
@@ 4.2 Ethical issues
Personification of the system Queries assuming the system as a human, and/or asking the system to perform tasks beyond its capabilities such as "Act like my boy-/girl-friend" and "Could you climb?". This category falls into a sensitive area depending on the purpose and definition of the service. In our case, queries regarding the pre-defined capabilities of the system do not have to be considered sensitive to prevent potential issues.
@@ 4.3 Service-sensitive issues
Error-inducing Instances of queries that elicit inaccurate or unexpected responses, such as hallucination and prompt injection, need to be addressed. For instance, media reports have highlighted cases where LLMs have provided affirmative responses to nonsensical questions, such as "Tell me the date when Samsung launched the latest iPhone?"
4.4 Considerations behind Taxonomy
Bias and Discrimination
Given our service’s primary focus on South Korea, sensitivity is inherently tied to cultural context. Labeling queries as sensitive (e.g., discrimination and controversial factuality) without considering nuanced cultural perspectives risks limiting free speech. However, while absolute neutrality is elusive, allowing potentially offensive queries could inadvertently convey corporate endorsement of sensitive viewpoints. To mitigate this, we have opted to avoid generating responses for such queries. Instead, we manage a whitelist, a component of our sensitive query classifier, to address excessive censorship.
Due to the nature mentioned above, the system focusing on a single cultural perspective could marginalize or incorrectly handle queries from users in different cultures. We acknowledge its importance but expanding this system to other countries would be a significant challenge.
Generality of Taxonomy
Most parts of the taxonomy and their insights discussed in Section 4 are not necessarily unique to the sociocultural contexts of South Korea but applicable to those of other countries. Consequently, we hope that the taxonomy serves as a reference for building similar services and as a benchmark framework for cross-cultural comparisons.
@@ 5 Analysis
We analyzed all user queries collected from 70 days since the launch of our service101010The distribution of queries collected after 70 days is presented in the Appendix. Instead of disclosing the exact numbers of active users and total queries due to confidentiality reasons, we provide a relative overview of the demographic distribution and the daily query volume in relation to the maximum number of daily queries, according to our taxonomy.
To effectively analyze the large volume of user queries, we developed a sensitive query classifier, which is further detailed in Section 5.2. This model enables us to examine the distribution of sensitive queries as discussed in Section 5.4. We use the taxonomy proposed in Section 4 for the distribution analysis. We also investigate how the distribution of sensitive queries changes in response to specific social issues that arose in news and social media.
@@ 5.1 User Demographics
5.2 Sensitive Query Classifier
5.2.1 Method
Model
We employ HyperCLOVA X backbone model Yoo et al. (2024), which is a state-of-the-art LLM for various benchmarks (including AI safety) in Korean. To account for the limited training data for safety available, we add a linear layer at the end of the backbone model to predict scores for 12 sensitive categories (with an additional safe class). Only the weights for the last layer are fine-tuned and the output of the last layer is used as a prefix in response generation.
We have constructed a dataset comprising 6,761 instances, combining publicly available Korean data Moon et al. (2020) and internal data generated by our human annotators. The detailed process of annotation will be illustrated in Section 5.3.
The performance is measured by a sampled test dataset. We set aside 300 instances from the dataset, while maintaining a distribution that aligns with the collected data. The accuracy of category classification is 85.3%111111We believe the performance is plausible for distribution analysis when considering the number of sensitive category and the large volume of input queries. Also, see the real-world performance presented in the end of this Section., and 157 out of 175 queries (89.7%) are correctly predicted as safe. The overall accuracy is thus 87% on the internal evaluation.
Rule-based Adjustment
The rule-based adjustment module aims to mitigate unexpected classifier behaviors in the service environment. By employing sentence-level regular expressions, the module associates queries with desired responses when the classifier either over-blocks safe queries or fails to detect sensitive ones. These rules, consisting of whitelists to make a response for over-blocked queries and blacklists to block for sensitive queries, guide the adjustment process. When a query requires adjustment, the model generates a response using the input query prefixed with the appropriate category (safe or sensitive).
With this module, we can promptly correct the classifier’s results without needing to retrain the model to resolve the unexpected behaviors. Additionally, the rules have been incorporated into the training data for future versions of the classifier.
Offline Test and Feedback Loop
To ensure the reliability of the classifier, we conduct offline tests due to potential disparities between the distribution of collected data and the real user inputs. Each day during the service, we sampled around 50 instances classified into sensitive categories, and had labelers to assess the accuracy of classification results121212We prioritize more conservative blocking rather than generating a response to a harmful query.. We have defined four categories: MustSafe, LookSafe, Harm, and CannotDecide. We have evaluated 3,692 examples, with 589 classified as MustSafe, 156 as LookSafe, 2,201 as Harm, and 746 as CannotDecide. The overall precision for the harm class is calculated at 74.7. Initially, the precision stood at 67.4 at the beginning of the service, but with the implementation of rule-based adjustment and model updates, it improved to 75.2 in the most recent offline test. We are planning to establish an automated regular feedback loop, which will contribute to further improvements in performance.
5.3 Considerations behind Method
Data Collection and Privacy
Our service terms explicitly state that user logs and input/output data are collected for purposes such as system improvement, statistical analysis, customer support (including error reporting and feedback analysis), and monitoring. This data is retained for up to nine months. Users who object to their data being used for research or AI development can request that it not be utilized. All user identifiers within the logs are encrypted and excluded from training and analysis processes.
Human Annotations
Our annotators were experienced full-time linguistics specialists131313Due to the employment type, their hard work on this project did not involve additional money incentives. with extensive backgrounds in NLP, ranging from 10 to 30 years. While their expertise was invaluable, they were not specifically trained to handle sensitive content. To minimize potential distress, we prioritized leveraging the inherent capabilities of HyperCLOVA X backbone model. Aligning with the model’s strong performance on AI safety benchmarks and the partial overlap between its ethics principles (HyperCLOVA X Ethics Principles Yoo et al. (2024)) and our taxonomy, we anticipated the model’s ability to filter out common or overtly sensitive queries.
However, to ensure the capture of subtle nuances and optimal model alignment with our service, human annotation remained essential. Annotators were asked to label publicly available Korean data using our taxonomy, categorizing queries as sensitive, over-blocked, or ambiguous. To mitigate workload and maintain their well-being, we limited annotations to 50 instances per annotator per day and provided a skip option for ambiguous cases. Additionally, we emphasized the importance of immediate cessation and reporting of any discomfort or mental fatigue. Annotators had the option to opt out of the task entirely, and we minimized their overall workload to alleviate stress.
System Transparency and Misclassification
While we recognize the potential risks associated with over-blocking or under-blocking queries, achieving 100% accuracy in content moderation is inherently challenging. To mitigate these risks, we prioritize user safety by explaining the reason of blocking and implementing a feedback mechanism. Users can report misclassifications, allowing us to refine our rules and improve system performance. Even though we explicitly address these limitations in our Terms of Service, we believe that proactively preventing users from exposure to harmful content is essential.
[…]
마약(drug),사이트(site),사건(event),번호(number),범죄(crime),
연예인(celebrity),영화(movie),남자(male),의심(doubt) |
Age-restricted contents | 4.9 | 11.6 | 사진(image),여자(girl),섹스(sex),친구(friend),남자(boy) |
Privacy | 1.9 | 5.1 |
번호(number),주소(address),정보(information),집(home),마약(drug),비밀(secret),아이디(ID),
개인(private),{CelebrityName},등록(register),사실(fact),루머(rumor), 연예(entertainment),주가(stock price) |
Copyright infringement | 4.3 | 10.7 | 사이트(site),무료(free),사진(image),다운로드(download),소설(novel),영화(movie) |
Discrimination | 36.1 | 45.5 | 남자(male),여자(female),이유(reason),친구(friend),문제(problem) |
Suicide&self-harm | 1.6 | 17.2 | 자살(suicide),고통(pain),우울(depression),엄마(mom),죽음(death) |
Profanity | 2.4 | 5.3 | 욕/비속어(swear),단어(a word),뜻(meaning),표현(expression),친구(friend),사용(use),반말(talking down) |
Personification of system | 12.2 | 18.2 |
친구(friend),사진(image),여자(girl),이름(name),대화(conversation),대통령(president),
거짓말(lie),좌파(left-winger) |
High-stakes domains | <0.1 | 0.4 |
학부모(school parent),민원(civil complaint),표(ticket),사례(example),
위치(location),악성(viciousness),극단(extreme) |
Future prediction | 7.9 | 15.8 |
주식(stock),가능(possibility),투자(investment),전망(prediction),주가(stock price),가격(price),
시장(market),비트(bit),공매도(short selling),금지(prohibition),변화(change) |
Controversial factuality | 17.8 | 32.3 |
대통령(president),땅(country),나라(nation),문제(problem),법(law),이유(reason),정치(politic),
배아(embryo),윤리(ethics),생명(life),길고양이(stray cat),세포(cell) |
Error-inducing | 1.1 | 3.8 |
꿈(dream),삼촌(uncle),인간/인류(human),인공지능(AI),멸망(doom),
지배(domination),세계(world),지구(earth) |
[…]